OpenStudy (anonymous):

How do I write a direct variation equation? Y=10 when x=2

OpenStudy (science0229):

The basic formula for direct variation is y=kx (k=constant), right?

OpenStudy (science0229):

So, 10=k*2 -------> k=5 So, the equation is y=5x

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

ok a direct variation equation would be written like this on a graph \[y=kx ,y=mx , m=slope , k=constant\]

OpenStudy (anonymous):

It says suppose y directly varies with x. Write a direct variation equation that relates x and y. Then find the value of y when x=8